UT Sep 2 (First Night)


OSA: DG.Huerta

ISp: G. Bechetti

ISA: B. Rothberg, O. Kuhn (Tucson)

AO: D. Miller (Tucson)

Mtn Manager: P. Hartley

Accomplished: Co-Pointing on LBCs

LBCs powered on OK - test biases look OK

MODS-2B with fiber swap worked, but when fibers were put back in place in the computer room, same issue appeared on readout - could be fiber is flaky?

Weather looks not promising - clouds overhead, virga visible, rain on all-sky earlier (before sunset)

UT 01:38 - power bump

UT 01:44 - AdSec failure on DX - could be due to power bump - D. Miller restarting

UT 01:58 - still closed, P. Hartley and G. Bechetti attending to power bump issues

UT 02:12 - AGW2 connected and both on/off axis are working

UT 02:34 - still closed - D. Miller reports DX AdSec is ready for SL use

UT 03:01 - still closed - LUCI temps rising - P. Hartley changed ICS offset to -4C making ICS temp 17C

UT 03:20 - DX primary Unknown state - troubleshooting

UT 03:33 - still closed, troubleshooting DX primary, re-swapped fibers on MODS2, now MODS2R shows bad readout (whereas when the swap occurred earlier both channels were OK).

UT 03:39 - PMC sub-system not running on DX

UT 04:06 - still closed - looks like Unknown error on DX is gone

UT 04:15 - LBC-red i-sloan images show a "drip" (looks like a tree branch) on chip 4 (closed dome testing), also on V-Bessel, I-Bessel, and R-Bessel, must be on the lenses (PA 90)

-will change PA to 0 to see if it rotates - next images show it on Chip4 but not as high on the chip

UT 04:27 - still closed - thermal PMC issue - D. Cox and S. Hooper working on issue

UT 04:31 - rotating LBC-R back to PA=90 to see if "twig" feature changes again

-went back to original position

-next try moving in elevation to see if this is something in the enclosure (can't for now)

-D. Thompson running grating checks with LUCI-1

-O. Kuhn working on LBC tracker tests - some examples of striping in tek chip images

UT 05:22 - still closed

UT 05:45 - PMC problem resolved

UT 06:10 - MODS-1 awake - sieve snaps taken, however, MODS-1 blue seems stuck after this, the quad cell counts are pegged at 0 and, consequently, imcslock fails on the blue side.

UT 06:18 - B. Rothberg restarted MODS1 IMCS - didn't seem to fix problem.

UT 06:25 - B. Rothberg quit and restarted the IC program. Also restarted the MODS1 IMCS service. Now we are seeing quad cell counts.

UT 06:54 - still closed

UT 08:07 - ECS - instrument alarm went off - LUCI-1 temps

-D. Thompson took image of new cube in park position - shows significant improvement

UT 08:13 - going to test if "twig" feature is visible in LBC-R at elev=60

UT 08:33 - seems to be gone

UT 08:47 - PSFGUI shows an error on the SX secondary, temp too high?

UT 09:00 - per G. Bechetti's request vent doors opened partially to allow chamber to cool

UT 09:51 - will attempt to open enclosure

-notified John Hill of our impending on-sky shenanigans

UT 10:01 - having trouble opening enclosure

UT 10:11 - trying to open again

UT 10:16 - moving to co-pointing field for collimation - 01:30:00 +50:00:00

UT 10:19 - running dohybrid

UT 10:29 - Dohybrid having difficulty finding pupils - rejecting pupils with 3000 pixels and 7 pixels, O. Kuhn stopping with Shift S - process hung but LBC still taking exposures

UT 10:31 - dohybrid finally quit, trying again

-WRS part of dohybrid is not fitting the coma properly, it thinks the center of the pupil is the bright spot

-glass thermals is 17.1/15.9C chamber is 14.9C outside is 12.7C

UT 10:37 - clouds appearing DG.Huerta closing enclosure

UT 10:47 - closed again

UT 11:31 - attempting to open again

UT 11:36 - same target as before, clearing out active optics

-dohybrid not removing coma

-cleared SX active optics, manually added +3000 z7 on SX

-added -2000 z8 on SX

-manually adding +2000 z11 to DX to try and recover central hole in pupil

-LBC-R looks better, but LBC-B got worse

UT 12:04 -->exiting dohybrid, going to run co-pointing script

-dohybrid is out of sequence,still taking exposure even though script stopped

-co-pointing image shows star which are round

-executed lbcrangebal

UT12:10 took a pair of images with the same copointing script.

-LBC-B psf looks triangular

UT 12:14 -going to take a pair of images with PA=180

UT 12:18 - skyflat sloan g and sloan i PA=0

UT 12:25 - seems like exposures are being taken out of sync between red and blue

sloan g ~20k, sloan i ~50k -- setting to scale = 0.6

UT 12:27 - LBC-B too few downloaded pixels on first exposure, LBC-R saturated first exposure

UT 12:29 - stopped at 4/5 for each

-attempting Uspec and z-sloan

-Saturated

UT 12:32 - attempting Uspec and z-sloan with scale=0.5

-first Uspec is ~50k

UT 12:34 - last Uspec shows too few downloaded pixels and jumbled image

UT 12:35 - closing enclosure

-powering off LBCs, MODS-1 going to sleep

MODS-2 fibers are still swapped


PROPOSED PLAN OF ACTION:
  • Telescope Wake-up ~0.5hr (test presets, TCS/PCS testing)
  • LBC Wake-up
    • Collimation - DoHybrid, then DOFPIA (~1hr)
    • Superfocus Sequence, PA=0, 180 (1hr)
    • Rotator Center Checks (pairs of 0,180; 90, 270) ~0.5hr
    • Test Science OB ~ 1hr
      • Sidereal
      • Non-Sidereal
    • Photometric Fields and Spectrophotometric Stars (to establish ZPs) ~ 1-1.5hrs
      • Requires multiple airmass visits (clear conditions, any seeing)
      • Cycle through all filters
      • This is now a high priority because we need to define ZPs
    • Twilight Flats
  • MODS On-sky checkout
    • Confirm centering of AGWs and guiding (test preset)
    • Basic Checkout
      • Rotator Centers
        • Collimate and then send track presets to the same field at PA=0, 180 and PA=90, 270. Analyze data with modsrc . modsrc works on both MODS and LBC images and it computes rotator centers for pairs of images which do not have to be taken at diametrically opposed position angles but can be, say, 90 deg apart instead.
      • Rotator Zero Points
        • Acquire 2 MOS masks and note the delta theta offset needed to align the mask. Ideally, the rotator zero point will be set (in PCSInstrument.conf) so that the mask or longslit is oriented at PA=0. The delta theta by which the instrument needs to be rotated to align the mask and field is the error in the rotator zeropoint. If the rotator zeropoint is updated, then a new transform needs to be determined. A 0.7 deg error in the rotator zeropoint translates to 0.73" at a distance 60" from the slit center. (If one of two stars separated by 60" is perfectly centered in a long-slit slit, the other will be on one edge of the slit unless the field is rotated by 0.7 degrees; since MOS mask alignment computes a rotational offset as well as a shift, the error in teh rotator zeropoint is corrected during a MOS alignment).
      • Transform:At minimum, check the existing transform by making a series of DETXY offsets with large (60") throw to compare orientation of deltaX, deltaY with slit.
      • Observe a standard star in all modes (30-45 min for 3 images). If clear, obtain the full set of 3 images per mode; if not, make a note in the object name and obtain only 1 image per mode.
      • Dice-5 imaging offsets
      • low priority check on orientation and scale of DETXY offsets, repeatability of these and also an opportunity to gather data on image quality.
    • Test Science OB
      • Sidereal
      • Non-Sidereal

-- BarryRothberg - 01 Sep 2019
Topic revision: r6 - 02 Sep 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback